A source of light is located at corner inside a 3m deep pool. If you were standing 4m away on the side of the pool. (1) will the light be totally reflected? (2) if you were standing directly above the source of light ,what will be the apparent depth of the light?
1
Expert's answer
2020-02-14T09:36:22-0500

The position of the light source =3m deep

Standing position on the side of the pool=4m

As per the question, the refractive index of the light is not given,

Let the refractive index of the light is "=\mu_1"

the refractive index of the air medium "\mu_2=1"

So

a)

"\dfrac{\sin (i)}{\sin (r)}=\dfrac{\mu_2}{\mu_1}"

"\sin (r)=\dfrac{\mu_2 \sin (i)}{\mu_1}"

"\sin (r)=\dfrac{ 3}{5\mu_1}"

if r=90

Then "\mu_1=\dfrac{3}{5}"


b)

Apperent depth ="d\dfrac{\mu_2}{\mu_1}"

"=3\dfrac{1}{\dfrac{3}{5}}=5cm"

So, apperent depth = 5cm